OpenStudy (anonymous):

What are all the real zeroes y = (x –12)^3 – 7? If an explanation could be provided too it'd be appreciated.

OpenStudy (anonymous):

zeros mean when y = 0

OpenStudy (anonymous):

Simplify the function first. It'll be easier.

OpenStudy (anonymous):

so set y = 0 meaning 0 = (x^2 - 24x + 144) - 7

OpenStudy (anonymous):

wooops it's cube

OpenStudy (anonymous):

set y =0. 0 = (x-12)^3 - 7 => (x-12)^3 = 7 => x-12 = cube rt (7) => x =[ cube rt (7) ]+12

OpenStudy (anonymous):

set y=0 then (x-12)^3=7, this gives x=12+cuberoot(7)=13.91(approx).